40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Apr 13 00:56:07 2016, gautam, Update, endtable upgrade, X endtable repopulation 
    Reply  Wed Apr 13 11:00:28 2016, Steve, Update, endtable upgrade, X endtable repopulation 
       Reply  Wed Apr 13 18:25:07 2016, gautam, Update, endtable upgrade, Lightwave health check 
          Reply  Fri Apr 15 18:38:12 2016, gautam, Update, endtable upgrade, Lightwave health check - NO IMPROVEMENT 
             Reply  Tue Apr 19 18:37:29 2016, gautam, Update, endtable upgrade, Laser swap + optical layout IR_modematch_19April2016_2.pdf
                Reply  Thu Apr 21 14:25:52 2016, gautam, Update, endtable upgrade, Green light recovered IMG_6567.JPGIMG_6568.JPGCVI_reflectivity.jpeg
                Reply  Tue Apr 26 15:22:35 2016, Steve, Update, endtable upgrade, Cleaning ETMX vacuum dirty window ETMX-Tvp.jpgETMX-TvpDetail.jpg
                   Reply  Tue Apr 26 23:19:42 2016, gautam, Update, endtable upgrade, Green aligned to arm - high order mode flashes seen IMG_2219.JPGIMG_2220.JPG
                      Reply  Thu Apr 28 00:41:08 2016, gautam, Update, endtable upgrade, more progress - Transmon PD installed 
                         Reply  Fri Apr 29 00:55:46 2016, gautam, Update, endtable upgrade, green PDH locked to Xarm 
                            Reply  Mon May 2 19:14:18 2016, gautam, Update, endtable upgrade, Optical layout almost complete ETMX_wedge.pdf
                               Reply  Thu May 5 03:05:37 2016, gautam, Update, endtable upgrade, ALS status update ALS_noiseSpec_5May2016_2.pdfCoherence_5May2016.pdfimage.jpeg
                                  Reply  Thu May 5 04:05:03 2016, ericq, Update, LSC, Aux X PDH checks nom_noise.pngnom_olg.png
                                     Reply  Thu May 5 14:03:52 2016, ericq, Update, LSC, Further Aux X PDH tweaks 2016-05-05_newals.pdf
                                  Reply  Thu May 5 14:05:01 2016, rana, Update, endtable upgrade, ALS status update 
                                  Reply  Thu May 5 21:28:44 2016, gautam, Update, endtable upgrade, Innolight PZT capacitance 
                               Reply  Thu May 19 16:29:20 2016, Steve, Update, endtable upgrade, Optical layout almost complete ETMX_4x3_closed.jpgsealedETMXenclosure.jpg
                                  Reply  Fri Sep 30 19:53:07 2016, gautam, Update, endtable upgrade, X end IR pickoff fiber coupled 
                                     Reply  Wed Oct 5 16:28:10 2016, gautam, Update, endtable upgrade, EX laser power monitor PD installed 
                   Reply  Wed Apr 27 15:04:47 2016, Steve, Update, endtable upgrade, Cleaning ETMX vacuum dirty window 1cETMX-Tcmp.jpg
                      Reply  Fri Apr 29 16:05:23 2016, gautam, Update, endtable upgrade, Cleaning ETMX vacuum dirty window IMG_0755.JPG
Message ID: 12075     Entry time: Wed Apr 13 18:25:07 2016     In reply to: 12074     Reply to this: 12079
Author: gautam 
Type: Update 
Category: endtable upgrade 
Subject: Lightwave health check 

[Koji,gautam]

Lightwave NPRO information:

Model: 126-1064-700

Serial Number: 337

Manufactured: December 1998!!

Details of checks performed:

Koji tuned the parameters on the laser controller and we observed the following:

  1. Turning "ADJ" to +10 and the pumping current all the way up to the maximum (2.62A) allowed us to recover an output power of 300mW, at a laser crystal temperature of ~45degrees
  2. The output power increased almost monotonically as a function of the laser crystal temperature - why? We were able to see powers as high as 250mW (at ADJ=0) for the maximum crystal temperature of ~60 degrees.
  3. We checked that we could believe the readout of the power meter by measuring the power using the Scientech power meter - we saw ~270mW after the Faraday with this meter, accounting for ~10% loss through the Faraday, this corresponds to an output power of 300mW (all this was done at ADJ=+10, DC=2.62A). I suspect that the display is dodgy though, because changing the Diode Current from 2.52A to 2.62A increased the output power by almost 100mW, which seems hard to believe?
  4. The Lightwave NPRO does not have heat dissipation fins attached - could this be affecting the power output somehow? In any case, this has to be rectified. So if we decide to keep the Lightwave NPRO, the layout will still need minor changes to accommodate the heat fins. Steve, do we have these in hand?

Way forward

Ericq has begun the characterization of the repaired Innolight. We checked that it outputs 1W of power. We will now have to perform the following measurements:

  • Frequency noise using PLL
  • AM/PM response of the PZT
  • Laser power output as a function of diode current - this will be useful for diagnostic purposes in the future
  • AUX temperature vs PSL temperature at which beatnotes can be found
  • Waist measurement - the mode matching and optical layout upstream of the doubling oven at least will have to be modified significantly

All of these will have to be done before installing this laser at the endtable.

I believe the consensus as of now is to go ahead with carrying out the above measurements. Meanwhile, we will keep the Lightwave NPRO on and see if there is some miraculous improvement. So the decision as to whether to use the Innolight is deferred for a day or two.

ELOG V3.1.3-